OpenStudy (anonymous):

Which point lies on a circle that is centered at A(-3, 2) and passes through B(1, 3)?

OpenStudy (jdoe0001):

if the circle's CENTER is at (-3,2) any point on the circle will be at a distance of "radius" from the center, that is 

OpenStudy (jdoe0001):

so just find the distance between A and B to get the radius and compare with other points, the distance must be the same for any other point on the circle \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -3}}\quad ,&{\color{blue}{ 2}})\quad B&({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

thank you jdoe

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

do you know what point the circle lies on

OpenStudy (jdoe0001):

hhmmm what did you get for the distance anyway?

OpenStudy (anonymous):

I got 2.25

OpenStudy (jdoe0001):

hmmm \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -3}}\quad ,&{\color{blue}{ 2}})\quad B&({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}}) \end{array}\qquad d = \sqrt{({\color{red}{ 1}}-{\color{red}{ -3}})^2 + ({\color{blue}{ 3}}-{\color{blue}{ 2}})^2} \ne 2.25\)

OpenStudy (jdoe0001):

notice that (1 - (-3)) => 1+3